# Hierarchical Position Embedding of Graphs with Landmarks and Clustering for Link Prediction

### v1

@@ 2.6. Design Insights from Theory
The key design questions to be addressed by our analysis are: what kind of, and how many, landmarks should be selected?
Landmark Selection and Graph Clustering. Theorem 3 states that, one should choose landmarks with sufficiently large $h$, e.g., high-degree nodes. However, such high-degree nodes may not always provide short paths for real-world graphs. For example, suppose all the hubs are located at one end of the network. The nodes at the other end of the network have to make a long detour via landmarks, even in order to reach nodes in local neighborhoods.
In order to better capture local graph structures and evenly distribute landmarks over the graph, we propose to partition the graph into clusters such that the nodes within a cluster tend to be densely connected, i.e., close to one another. Then we pick the node with the highest degree within each cluster as the landmark, as suggested by the analysis. Each landmark represents the associated cluster, and the distance between nodes can be captured by distance information associated with the cluster landmarks. We empirically find that such combination of clustering and landmark selection yields improved results.
Number of Landmarks. Although large number of landmarks $K(N)$ appears preferable, our analysis show that $K(N)$ does not drastically reduce distances, unless $K(N)$ is very large, e.g., $N^{1-\epsilon}$. A large number of landmarks can hamper scalability. Thus, we use only a moderate number of landmarks, and empirically find that setting $K(N)=O(\log N)$ suffices to yield good results.
@@ 3.1. Graph Clustering and Landmark Selection
For each cluster, we select the node with the highest degree as the landmark, where $\lambda_{k}$ denotes the landmark of cluster $C_{k}$. For node $v$, we compute the distances to landmarks $\lambda_{1},\cdots,\lambda_{K}$ to yield distance vector $D(v)$. Since $G$ may be disconnected, we extend the definition of distance $d(v,\lambda)$ such that, in case there exists no path from $v$ to $\lambda$, $d(v,\lambda)$ is set to $\bar{d}+1$ where $\bar{d}$ is the maximum diameter among the connected components. As mentioned earlier, we set number of clusters $K=\eta\log N$ where integer $\eta$ is a hyperparameter.
@@
A supporting argument for incorporating clustering into analysis can be made for BA networks. Theorem 3 states that, choosing $K(N)=O(\log N)$ landmarks from top-$(\log N)\cdot K(N)=O(\log^{2}N)$ nodes in degrees is optimal. We show that, $\log N$ landmarks chosen after FluidC clustering are indeed within top-$(\log N)^{2}$ degree centrality through simulation. Table 1 compares the rank of degrees of landmarks selected through FluidC clustering versus the rank of top-$(\log N)^{2}$ degree nodes in BA networks. We observe that the landmarks selected after clustering have sufficiently large degrees for optimality, i.e., all of their degrees are within top-$(\log N)^{2}$.
@@
One can expect similar effects from FluidC clustering for real-world graphs. Given the small number $(O(\log N))$ of clusters as input, the algorithm will detect locally dense clusters resulting in landmarks which have high degrees and are evenly distributed over the graph. Thus, landmark selection after clustering can be a good heuristic motivated by the theory.
@@
Table 1. Rank of degrees of landmark nodes in BA networks. For example, if $N=500$, all the landmarks selected from clustering are within the top 5.02%-degree nodes.
@@ 3.2. Membership Encoding with Graph Laplacian
For the nodes within the same cluster, we augment the nodes’ embeddings with information identifying that they are members to the same community, which we call membership. The membership is extracted from landmarks, and is based on the relative positional information among landmarks. Thus, not only the nodes within the same cluster have the same membership, but also the nodes of neighboring clusters have “similar” membership. The membership of a node is encoded into a membership vector (MV) which is combined with DV in computing the node embeddings.
We consider the graph consisting only of landmark nodes, and use graph Laplacian (Belkin and Niyogi, 2003) to encode their relative positions. A complete graph of landmarks is constructed where the edge weight between landmarks $u$ and $v$ is set to $e_{uv}=\exp(-d(u,v)^{2}/T)$ where $T$ denotes normalizing parameter of heat kernel. Let $\hat{A}\in\mathbb{R}^{K\times K}$ denote the weighted adjacency matrix. The normalized graph Laplacian is given by $L=I-\Delta^{-\frac{1}{2}}\hat{{A}}\Delta^{-\frac{1}{2}}$ where degree matrix $\Delta$ is given by $\Delta_{ii}=\sum_{j}\hat{A}_{ij}$. The eigenvectors of $L$ are used as MVs, i.e., for node $v\in C_{k}$, the MV of node $v$, denoted by $M(v)$, is the eigenvector of $L$ associated with landmark $\lambda_{k}$. MV provides positional information in addition to DV, using the graph of upper hierarchy, i.e., landmarks. We use a random flipping of the signs of eigenvectors to resolve ambiguity (Dwivedi et al., 2020). A problem with Laplacian encoding is the time complexity, of which the eigendecomposition is cubic in network size. Previous approaches used a subset of eigenspectrum (Dwivedi et al., 2020; Kreuzer et al., 2021). However, the landmark graph has small size, i.e., $K=O(\log N)$. Thus, our model can exploit the full spectrum, enabling accurate representation of landmark graphs.
@@ 3.3. Cluster-group Encoding
We propose cluster-group encoding as an additional model optimization.
Several neighboring clusters are further grouped into a macro-cluster. The embeddings of nodes in a macro-cluster use an encoder specific to that macro-cluster. The motivation is that, using a separate encoder per local region may facilitate capturing attributes specific to local structures or latent features of the community, which is important for link prediction.
@@
where $x_{v}$ is the input node feature, $M(v)$ and $D(v)$ are MV and DV, $\phi$ is membership/distance encoder, and $\oplus$ denotes concatenation. $f_{i(\cdot)}$ denotes the encoder specific to macro-cluster $\Phi_{i(\cdot)}$. We use MLPs for $\phi(\cdot)$ and $\Phi_{i(\cdot)}$. The number of macro-cluster, denoted by $R$, is set such that $K$ is a multiple of $R$. Each macro-cluster contains $K/R$ clusters. Cluster-group encoding requires $R$ encoders, one per macro-cluster. To limit the model size, we empirically set $R=\min(15,\lfloor K/\eta\rfloor)$. For ease of implementation, $G$ is first partitioned into $R$ macro-clusters, then each macro-cluster is partitioned into $K/R$ clusters. This is simpler than, say, partitioning $G$ into $K$ (smaller) clusters and grouping them back to macro-clusters. Next, embedding $z_{v}$ is input to GNN layers given by
$h^{l}_{v}=\textrm{GNN}(h^{l-1}_{v},\mathbf{A}),\quad\textrm{for }l=1,2,\cdots$
@@ 4. Property of HPLC as Node embedding
It is argued in (Srinivasan and Ribeiro, 2020) that positional node embeddings are good at link prediction tasks as follows. The positional embeddings preserve the relative positions of nodes, which enables differentiating isomorphic nodes according to their positions and identifying the closeness between nodes. We claim that HPLC qualifies as positional node embeddings according to Definition 1 as follows.
In HPLC, landmark-based distance function, eigenvectors of graph Laplacian with random flipping, and graph clustering method are $\mathcal{G}$-equivariant functions of $\mathbf{A}$ ignoring the node features $X$. The landmark-based distance vectors and its functions clearly satisfy the equivariance property. The eigenvectors of graph Laplacian are permutationally equivariant under the node permutation, i.e., switching of corresponding rows/columns of adjacency matrix. Also, the Laplacian eigenmap can be regarded as a $\mathcal{G}$-equivariant function in the sense of expectation, if it is combined with random sign flipping of eigenvectors (Srinivasan and Ribeiro, 2020). In case of multiplicity of eigenvalues, we can slightly perturb the edge weights of landmark graphs to obtain simple eigenvalues (Poignard et al., 2018). The graph clustering in HPLC is a randomized method, because initially $K$ central nodes are selected at random. Thus, embedding output $Z$ of HPLC is a function of $\mathbf{A},X$, and random noise, which proves our claim that HPLC qualifies as a positional node embedding. Moreover, HPLC is more expressive than standard message-passing GNNs as follows. HPLC is trained to predict the link between a node pair based on their positional embeddings. The embeddings are learned over the joint distribution of distances from node pairs to common landmarks which are spread out globally over the graph. By contrast, traditional GNNs learn the embeddings based on the marginal distributions of local neighborhoods of node pairs. By a similar argument to Sec. 5.2 of (You et al., 2019), HPLC induces higher mutual information between its embeddings and the target similarity (likelihood of link formation) of node pairs based on positions, and thus has higher expressive power as compared to standard GNNs.
@@ 5.2. Space Complexity
Overall, the time and space complexity of HPLC is low, and our experiments shows that HPLC handles large or dense graphs well.
@@
Table 2. Link prediction results on various datasets. All baselines and our method were evaluated for 10 repetitions. Bold denotes the best performance, and Italic indicates the second best performance. We used a single NVIDIA RTX 3090 with 24GB memory on all datasets except CITATION2 and A100 GPU with 40GB memory on CITATION2. - indicates ‘out-of-memory’ (OOM). Some baselines suffered from OOM on large graphs due to the high memory usage from storing a large number of shortest paths, attention weights, or aggregation of hidden embedding vectors, etc. Similar OOM results as well as poor performance of those baselines were reported in (Wang et al., 2022) and (Zhao et al., 2022). For SEAL and GCN+DE, we trained 2% of training data and evaluated 1% of both validation and test set respectively on CITATION2. We trained 15% of training data but evaluated all of the validation and test sets on COLLAB. Both implementations followed the guideline on the official GitHub of SEAL-OGB. ‘Avg.’ denotes the average of performance metrics, and ‘H.M’ indicates their harmonic mean. (-) means that we do not report the average and harmonic mean due to OOM.
@@ 6.1. Experimental setting
Datasets. Experiments were conducted on 7 datasets widely used for evaluating link prediction. For experiments on small graphs, we used PubMed, Cora, Citeseer, and Facebook. For experiments on dense or large graphs, we chose DDI, COLLAB, and CITATION2 provided by OGB (Hu et al., 2020). Detailed statistics and evaluation metrics associated with the datasets are provided in Table 10 in Appendix LABEL:dataset.
@@
Evaluation metrics. Link prediction was evaluated based on the ranking performance of positive edges in the test data over negative ones. For COLLAB and DDI, we ranked all positive and negative edges in the test data, and computed the ratio of positive edges which are ranked in top-$k$. We did not utilize validation edges for computing node embeddings when we predicted test edges on COLLAB. In CITATION2, we computed all positive and negative edges, and calculated the reverse of the mean rank of positive edges. Due to high complexity when evaluating SEAL, we only trained 2% of training set edges and evaluated 1% of validation and test set edges respectively, as recommended in the official GitHub of SEAL. For Cora, Citeseer, PubMed, and Facebook, we utilized Area Under ROC Curve (AUC). If applicable, we calculated the average and harmonic mean (HM) of the measurements. HM penalizes the model for very low scores, thus is a useful indicator of robustness.
Hyperparameters. We used GCN as our base GNN encoder. MLP is used in decoders, except GAE. All baselines do not use edge weights. During training, negative edges are randomly sampled at a ratio of 1:1 with positive edges. The details of hyperparameters are provided in Appendix E.
@@ 6.2. Results
Experimental results are summarized in Table 2. HPLC outperformed the baselines on most datasets. HPLC achieved large performance gains over GAE combined with GCN on all datasets, which are 88.9% on DDI, 27.0% on COLLAB, 12.7% on Citeseer, 7.1% on Cora, and 1.4% on CITATION2. HPLC showed superior performance over SEAL, achieving gains of 167% on DDI, and 12.0% on Citeseer. We compare HPLC with other distance-based methods. Compared to GCN+DE which encodes distances from a target node set whose representations are to be learned, or to P-GNN which uses distances to random anchor sets, HPLC achieved higher performance gains by a large margin. HPLC also outperformed other positional encoding methods such as GCN+LPE, Graph Transformer+LPE, and PEG-DW+. The results show that the estimate of inter-node distances via landmarks can be effective for representing the positional information of nodes.
SEAL and NBF-net performed poorly on DDI which is a highly dense graph. Since the nodes of DDI have a large number of neighbors, the enclosing subgraphs are both very dense and large, and the model struggles with learning the representations of local structures or paths between nodes. By contrast, HPLC achieved the best performance on DDI, demonstrating its effectiveness on densely connected graphs.
@@ 7.1. Model Components
Table 3 shows the performances in the ablation analysis for the model components. The components denoted by“DV”, “CE” and “MV” columns in Table 3 indicate the usage of distance vector, cluster-group encoding and membership vector, respectively. “DV” is the default component of HPLC. The results show that the hierarchical components of HPLC, i.e., ‘CE” and “MV”, indeed contribute to the performance improvement.

### v2

@@ 2.6. Design Insights from Theory
The key design questions we would like to address are: what kind of, and how many, landmarks should be selected?
Landmark Selection and Graph Clustering. Theorem 3 states that, one should choose landmarks with sufficiently large $h$, e.g., high-degree nodes. However, such high-degree nodes may not always provide short paths for real-world graphs. For example, suppose all the hubs are located at one end of the network. The nodes at the other end of the network need to make a long detour via landmarks, even in order to reach nodes in local neighborhoods.
In order to better capture local graph structure and evenly distribute landmarks over the graph, we propose to partition the graph into clusters such that the nodes within a cluster tend to be densely connected, i.e., close to one another. Next, we propose to pick the node with the highest degree within each cluster as the landmark, as suggested by the analysis. Each landmark represents the associated cluster, and the distance between nodes can be captured by distance vectors associated with the cluster landmarks. We empirically find that the combination of landmark selection and clustering yields improved results.
Number of Landmarks. Although large $K(N)$ appears preferable, our analysis show that $K(N)$ does not drastically reduce distances, unless $K(N)$ is very large, e.g., $N^{1-\epsilon}$. A large number of landmarks can hamper scalability. Thus, we use only a moderate number of landmarks as suggested by the analysis. Indeed, we empirically found that setting $K(N)=O(\log N)$ suffices to yield good results.
@@ 3.1. Graph Clustering and Landmark Selection
For each cluster, we select the node with the highest degree as the landmark, where $\lambda_{k}$ denotes the landmark of cluster $C_{k}$. For node $v$, we compute the distances to landmarks $\lambda_{1},\cdots,\lambda_{K}$ to yield distance vector $D(v)$. Since $G$ may be disconnected, we extend the definition of distance $d(v,\lambda)$ such that, in case there exists no path from $v$ to $\lambda$, $d(v,\lambda)$ is set to $\bar{d}+1$ where $\bar{d}$ is the maximum diameter among the connected components. As stated earlier, we set number of clusters $K=\eta\log N$ where integer $\eta$ is a hyperparameter.
@@
A supporting argument for incorporating clustering into analysis can be made for BA graphs. Theorem 3 states that, choosing $K(N)=O(\log N)$ landmarks from top-$(\log N)\cdot K(N)=O(\log^{2}N)$ nodes in degrees is optimal. We show that, $\log N$ landmarks chosen after FluidC clustering are indeed within top-$(\log N)^{2}$ degree centrality through simulation. Table 1 compares the rank of degrees of landmarks selected after FluidC clustering versus the rank of top-$(\log N)^{2}$ degree nodes in BA graphs. We observe that the landmarks selected after clustering have sufficiently large degrees for optimality, i.e., all of their degrees are within top-$(\log N)^{2}$.
@@
One can expect similar effects from FluidC clustering for real-world graphs. Given the small number $(O(\log N))$ of clusters as input, the algorithm will detect locally dense clusters resulting in landmarks which have high degrees and tend to be evenly distributed over the graph. Thus, landmark selection after clustering can be a good heuristic motivated by the theory.
@@
Table 1. Rank of degrees of landmark nodes in BA graphs. For example, if $N=500$, all the landmarks selected from clustering are within the top 5.02%-degree nodes.
@@ 3.2. Membership Encoding with Graph Laplacian
For the nodes within the same cluster, we augment the nodes’ embeddings with information identifying that they are members of the same community, which we call membership. The membership is extracted from landmarks, and is based on the relative positional information among landmarks. Thus, not only the nodes within the same cluster have the same membership, but also the nodes of neighboring clusters have “similar” membership. The membership of a node is encoded into a membership vector (MV) which is combined with DV in computing the node embeddings.
We consider the graph consisting only of landmark nodes, and use graph Laplacian (Belkin and Niyogi, 2003) to encode their relative positions as follows. The landmark graph uses the edge weight between landmarks $u$ and $v$ which is set to $e_{uv}=\exp(-d(u,v)^{2}/T)$ where $T$ denotes normalizing parameter of heat kernel. Let $\hat{A}\in\mathbb{R}^{K\times K}$ denote the weighted adjacency matrix. The normalized graph Laplacian is given by $L=I-\Delta^{-\frac{1}{2}}\hat{{A}}\Delta^{-\frac{1}{2}}$ where degree matrix $\Delta$ is given by $\Delta_{ii}=\sum_{j}\hat{A}_{ij}$. The eigenvectors of $L$ are used as MVs, i.e., for node $v\in C_{k}$, the MV of node $v$, denoted by $M(v)$, is the eigenvector of $L$ associated with landmark $\lambda_{k}$. MV provides positional information in addition to DV, using the graph of upper hierarchy, i.e., landmarks. We use a random flipping of the signs of eigenvectors to resolve ambiguity (Dwivedi et al., 2020). A problem with Laplacian encoding is the time complexity: one needs the eigendecomposition whose complexity is cubic in network size. Thus, previous approaches had limitations, i.e., used only a subset of eigenspectrum (Dwivedi et al., 2020; Kreuzer et al., 2021). However, in our method, the landmark graph has small size, i.e., $K=O(\log N)$. Thus, we can exploit the full eigenspectrum, enabling more accurate representation of landmark graphs.
@@ 3.3. Cluster-group Encoding
We propose cluster-group encoding as an additional model optimization as follows. Several neighboring clusters are further grouped into a macro-cluster. The embeddings of nodes in a macro-cluster use an encoder specific to that macro-cluster. The motivation is that, using a separate encoder per local region may facilitate capturing attributes specific to local structures or latent features of the community, which is important for link prediction.
@@
where $x_{v}$ is the input node feature, $M(v)$ and $D(v)$ are MV and DV, $\phi$ is membership/distance encoder, and $\oplus$ denotes concatenation. $f_{i(\cdot)}$ denotes the encoder specific to macro-cluster $\Phi_{i(\cdot)}$. MLPs are used for $\phi(\cdot)$ and $\Phi_{i(\cdot)}$. The number of macro-cluster, denoted by $R$, is set such that $K$ is a multiple of $R$. Each macro-cluster contains $K/R$ clusters. Cluster-group encoding requires $R$ encoders, one per macro-cluster. To limit the model size, we empirically set $R=\min(15,\lfloor K/\eta\rfloor)$. For ease of implementation, $G$ is first partitioned into $R$ macro-clusters, then each macro-cluster is partitioned into $K/R$ clusters. This is simpler than, say, partitioning $G$ into $K$ (smaller) clusters and grouping them back to macro-clusters.
Next, $z_{v}$ is input to GNN whose $l$-th layer output $h^{l}_{v}$ is given by
$h^{l}_{v}=\textrm{GNN}(h^{l-1}_{v},\mathbf{A}),\quad\textrm{for }l=1,2,\ldots,L$
@@ 4. Property of HPLC as Node embedding
It is argued in (Srinivasan and Ribeiro, 2020) that positional node embeddings are good at link prediction tasks: the positional embeddings preserve the relative positions of nodes, which enables differentiating isomorphic nodes according to their positions and identifying the closeness between nodes. We claim that HPLC qualifies as positional node embeddings according to Definition 1 as follows.
In HPLC, landmark-based distance function, eigenvectors of graph Laplacian with random flipping, and graph clustering method are $\mathcal{G}$-equivariant functions of $\mathbf{A}$ ignoring the node features $X$. The landmark-based distance vectors and its functions clearly satisfy the equivariance property. The eigenvectors of graph Laplacian are permutationally equivariant under the node permutation, i.e., switching of corresponding rows/columns of adjacency matrix. Also, the Laplacian eigenmap can be regarded as a $\mathcal{G}$-equivariant function in the sense of expectation, if it is combined with random sign flipping of eigenvectors (Srinivasan and Ribeiro, 2020). In case of multiplicity of eigenvalues, we can slightly perturb the edge weights of landmark graphs to obtain simple eigenvalues (Poignard et al., 2018). The graph clustering in HPLC is a randomized method, because initially $K$ central nodes are selected at random. Thus, embedding output $Z$ of HPLC is a function of $\mathbf{A},X$, and random noise, which proves our claim that HPLC qualifies as a positional node embedding.
Moreover, HPLC is more expressive than standard message-passing GNNs as follows. HPLC is trained to predict the link between a node pair based on their positional embeddings. The embeddings are learned over the joint distribution of distances from node pairs to common landmarks which are spread out globally over the graph. By contrast, traditional GNNs learn the embeddings based on the marginal distributions of local neighborhoods of node pairs. By a similar argument to Sec. 5.2 of (You et al., 2019), HPLC induces higher mutual information between its embeddings and the target similarity (likelihood of link formation) of node pairs based on positions, and thus has higher expressive power than standard GNNs.
@@ 5.2. Space Complexity
Overall, the time and space complexity of HPLC is low, and our experiments shows that HPLC handles large or dense graphs well. We also measured the actual resource usage, and the related results are provided in Appendix D.
@@
Table 2. Link prediction results on various datasets. All baselines and our method were evaluated for 10 repetitions. Bold denotes the best performance, and Italic indicates the second best performance. We used a single NVIDIA RTX 3090 with 24GB memory on all datasets except CITATION2 and A100 GPU with 40GB memory on CITATION2. - indicates ‘out-of-memory’ (OOM). Some baselines suffered from OOM on large graphs due to the high memory usage from storing a large number of shortest paths, attention weights, or aggregation of hidden embedding vectors, etc. Similar OOM results as well as poor performance of those baselines were reported in (Wang et al., 2022) and (Zhao et al., 2022). For SEAL and GCN+DE, we trained 2% of training data and evaluated 1% of both validation and test set respectively on CITATION2. We trained 15% of training data but evaluated all of the validation and test sets on COLLAB. Both implementations followed the guideline on the official GitHub of SEAL-OGB. ‘Avg.’ denotes the average of performance metrics, and ‘H.M’ indicates their harmonic mean. (-) in ‘Avg. (H.M.)’ means that we do not report the average and harmonic mean due to OOM.
@@ 6.1. Experimental setting
Datasets. Experiments were conducted on 7 datasets widely used for evaluating link prediction. For experiments on small graphs, we used PubMed, Cora, Citeseer, and Facebook. For experiments on dense or large graphs, we chose DDI, COLLAB, and CITATION2 provided by OGB (Hu et al., 2020). Detailed statistics and evaluation metrics of the datasets are provided in Table 9 in Appendix E.
@@
Evaluation metrics. Link prediction was evaluated based on the ranking performance of positive edges in the test data over negative ones. For COLLAB and DDI, we ranked all positive and negative edges in the test data, and computed the ratio of positive edges which are ranked in top-$k$. We did not utilize validation edges for computing node embeddings when we predicted test edges on COLLAB. In CITATION2, we computed all positive and negative edges, and calculated the reverse of the mean rank of positive edges. Due to high complexity when evaluating SEAL, we only trained 2% of training set edges and evaluated 1% of validation and test set edges respectively, as recommended in the official GitHub of SEAL. The metrics for DDI, COLLAB and CITATION2 are chosen according to the official OGB settings  (Hu et al., 2020). For Cora, Citeseer, PubMed, and Facebook, Area Under ROC Curve (AUC) is used as the metric similar to prior works (Kipf and Welling, 2017; Wang et al., 2022). If applicable, we calculated the average and harmonic mean (HM) of the measurements. HM penalizes the model for very low scores, thus is a useful indicator of robustness.
Hyperparameters. We used GCN as our base GNN encoder. MLP is used in decoders, except GAE. None of the baseline methods used edge weights. During training, negative edges are randomly sampled at a ratio of 1:1 with positive edges. The details of hyperparameters are provided in Appendix F.
@@ 6.2. Results
Experimental results are summarized in Table 2. HPLC outperformed the baselines on most datasets. HPLC achieved large performance gains over GAE combined with GCN on all datasets, which are 88.9% on DDI, 27.0% on COLLAB, 12.7% on Citeseer, 7.1% on Cora, and 1.4% on CITATION2. HPLC showed superior performance over SEAL, achieving gains of 167% on DDI, and 12.0% on Citeseer. We compare HPLC with other distance-based methods. Compared to GCN+DE which encodes distances from a target node set whose representations are to be learned, or to P-GNN which uses distances to random anchor sets, HPLC achieved higher performance gains by a large margin. HPLC outperformed other positional encoding methods, e.g., GCN+LPE, Graph Transformer+LPE, and PEG-DW+. The results show that the proposed landmark-based representation can be effective for estimating the positional information of nodes.
SEAL and NBF-net performed poorly on DDI which is a highly dense graph. Since the nodes of DDI have a large number of neighbors, the enclosing subgraphs are both very dense and large, and the models struggle with learning the representations of local structures or paths between nodes. By contrast, HPLC achieved the best performance on DDI, demonstrating its effectiveness on densely connected graphs.
@@ 7. Ablation study
In this section, we provide the ablation study. Additional ablation studies on graph clustering algorithms and node centrality are provided in Appendix C.
@@ 7.1. Model Components
Table 3 shows the performances in the ablation analysis for the model components. The components denoted by“DV”, “CE” and “MV” columns in Table 3 indicate the usage of distance vector, cluster-group encoding and membership vector, respectively. The results show that all of the hierarchical components, i.e., “DV”, “CE” and “MV”, indeed contribute to the performance improvement.